OpenStudy (anonymous):

Simplify the given expression: (1 + 2i) • (5 − 3i)

OpenStudy (anonymous):

Use foil.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

1 * 5 = a 2i * 5 = b 1 * -3i = c 2i * -3i = d a + b - c - d Simplify.

OpenStudy (anonymous):

i still dont understand math is a little hard for me to get

OpenStudy (yanasidlinskiy):

Jcpd910. Mind explaining?

OpenStudy (anonymous):

1 * 5 = a =1 2i * 5 = b =10i So you plug in the values into the equation a = 1 b = 10i 1 + 10i - c - d

OpenStudy (anonymous):

i get it thanks !

OpenStudy (anonymous):

No problem. :D

OpenStudy (anonymous):

it doesnt give me any of my choices .2 − 15i 3 + 2i 5 − 6i 11 + 7i

OpenStudy (anonymous):

unless im still doing it wrong

OpenStudy (anonymous):

Show me your work please.

OpenStudy (anonymous):

1 * 5 = a =1 2i * 5 = b =10i Okayy Here Why Did You Say 5* 1 = 1?

OpenStudy (anonymous):

Oh, I'm sorry I meant 5. lol

OpenStudy (anonymous):

thats okay so this is what i got 1*5=5 2i*5=10i 1 *-3i=-3i 2 * -3i=-6i

OpenStudy (anonymous):

Okay, I'm getting confused. Are you using i as an imaginary number or as a variable?

OpenStudy (anonymous):

imaginary number

OpenStudy (anonymous):

Okay, well let's see now. (1 + 2i) • (5 − 3i) 5 + 10i - 3i - 6i^2 Do you know what i^2 =?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Yes, so 5 + 10i - 3i -(6(-1) 5 + 10 - 3i + 6 11 + 7i

OpenStudy (anonymous):

Sorry, I thought i was a variable.

OpenStudy (anonymous):

So now you got your answer.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem. :)